OpenStudy (anonymous):

Need Help. write an equation of a parabola with vertex at the origin and the given focus (0 -5)

jimthompson5910 (jim_thompson5910):

The origin is (0,0)

jimthompson5910 (jim_thompson5910):

is the focus above or below the vertex?

OpenStudy (anonymous):

Above?

jimthompson5910 (jim_thompson5910):

Vertex: (0,0) Focus: (0,-5)

jimthompson5910 (jim_thompson5910):

is the focus above or below the vertex?

OpenStudy (anonymous):

Below?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what is the vertical distance from the focus to the vertex?

OpenStudy (anonymous):

-5 units? I'm not sure, lol not very good at match at all!

jimthompson5910 (jim_thompson5910):

distance is never negative

jimthompson5910 (jim_thompson5910):

it's actually +5 or just 5

jimthompson5910 (jim_thompson5910):

so that means the focal distance, p, is p = 5

OpenStudy (anonymous):

Oh ok. I'll remember that.

jimthompson5910 (jim_thompson5910):

the vertex is (h,k) = (0,0) so, h = 0 k = 0

jimthompson5910 (jim_thompson5910):

plug p = 5, h = 0, and k = 0 into 4p(y-k) = (x-h)^2 and solve for y. Tell me what you get.

OpenStudy (anonymous):

20y-20=x^2-0?

jimthompson5910 (jim_thompson5910):

k = 0, not k = 1

OpenStudy (anonymous):

oh duh.. so 20y-0=x^2-0?

jimthompson5910 (jim_thompson5910):

that's the same as 20y = x^2 then you can solve for y to get y = ???

OpenStudy (anonymous):

1/20x^2?? I have no idea... I'm drawing a blank.

jimthompson5910 (jim_thompson5910):

yes, correct now because the focus is below the vertex, this means the parabola opens downward so the leading coefficient is negative making the final answer to be \[\Large y = -\frac{1}{20}x^2\]

OpenStudy (anonymous):

Okay!! Thank you so much!! You helped a lot!

jimthompson5910 (jim_thompson5910):

you're welcome